Eliminate arbitrary constants from z=(x-a)2+(y-b)2 to form the partial differential equation.


1
Expert's answer
2021-07-15T05:32:18-0400
"z=(x-a)^2+(y-b)^2"

Differentiating partially with respect to "x" and "y," we get


"\dfrac{\partial z}{\partial x}=2(x-a)"

"\dfrac{\partial z}{\partial y}=2(y-b)"

Squaring and adding these equations, we have


"(\dfrac{\partial z}{\partial x})^2+(\dfrac{\partial z}{\partial y})^2=(2(x-a))^2+(2(y-b))^2"

"(\dfrac{\partial z}{\partial x})^2+(\dfrac{\partial z}{\partial y})^2=4((x-a)^2+(y-b)^2)"

Since "(x-a)^2+(y-b)^2=z," we get


"(\dfrac{\partial z}{\partial x})^2+(\dfrac{\partial z}{\partial y})^2=4z"